A rectangle has a perimiter of 90 cm the length of the rectangle is 25 cm more than its width what is the number of square centimeters in the area of the rectangle? please help me i am completely confused
Answer by rfer(16322)

find the sides first, then the area
P=2w+2l
2x+2(x+25)=90
2x+2x+50=90
4x=40
x=10cm
x+25=35cm
-----------------
A=wl
A=10*35
A=350 sq cm
-----------------
Hope that helps
Bob